Hi I need help on my math package any tutors willing to help me would be greatly appreciate it. I just moved and need help to learn the topics. Julia can clean an attic in 10.2 hours. Rob can clean the same attic in 14.6 hours. How long would it take for the two of them if the worked together? Do we add the two numbers and divide by 2 to get 12.4 is that right?
Answer by stanbon(75887)

Julia can clean an attic in 10.2 hours. Rob can clean the same attic in 14.6 hours. How long would it take for the two of them if the worked together?
-----
Julia rate: 1/10.2 job/hr
Rob rate: 1/14.6 job/hr
Together rate: 1/x job/hr
====================
Equation:
rate + rate = together rate
1/10.2 + 1/14.6 = 1/x
14.6x + 10.2x = 10.2*14.6
24.8x = 10.2*14.6
x = 6 hrs (time to do the job together)
========================
Cheers,
Stan H.
------------------------